If m varies directly to the square root of k, and k = 196 when m = 27. Find m when k = 4,900.
Answer by sachi(548)

If m varies directly to the square root of k,
or m=c sqrtk where c is the const of variation
so c=m/sqrtk=27/sqrt196=27/14 as k = 196 when m = 27.
find m when k = 4,900.
m=27/14*sqrt4900=(27/14)*70=135